OpenStudy (anonymous):

y=10x^2-15x+5 & y=5x+5 find the point of intersection of the wo systems

OpenStudy (matt101):

At the point of intersection, one pair of coordinates satisfies both equations. Therefore, you can set them equal to each other (since the y-values are the same) and solve for x: 10x^2 -15x + 5 = 5x + 5 10x^2 - 20x = 0 x^2 - 2x = 0 x(x-2) = 0 x = 0, 2 Sub the points into one of the original equations to find the points of intersection. y = 5(0) + 5 = 5 y = 5(2) +5 = 15 There are two points of intersection: (0,5) and (2,15).

OpenStudy (anonymous):

thanks very much